OpenStudy (anonymous):

PLEASE HELP! Given a soda can with a volume of 36 and a diameter of 4, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank). (4 points)

OpenStudy (anonymous):

In order to find the volume of the cone, you need to solve for the height of the soda can and use that height in the volume for the cone. I found the height for the can to be 2.86 (although that is not accurate at all!) and plugged in that height for the cone and got a volume of 11.97 units cubed.

OpenStudy (ddcamp):

If you have a cylinder, the volume of a cone that fits perfectly inside it is 1/3 the volume of the cylinder.

OpenStudy (anonymous):

Vcone=13∗Basearea∗height

OpenStudy (anonymous):

Vcyl=Basearea∗height

OpenStudy (anonymous):

36/3 is equal to 12

OpenStudy (ddcamp):

Formula for a Cylinder: \[V = \pi r^2 h\] For a cone: \[V = \frac{1}{3} \pi r^2 h\]

OpenStudy (ddcamp):

Since the radius and height of the cone and the cylinder will be the same (since the cone fits perfectly), we can use these equations to get: \[V_{cone} = \frac{1}{3}V_{cylinder}\]

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

so the answer is 12?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

where did yu get the 3 in 36/3